OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies inversely as x and y=11 when x=6.

OpenStudy (anonymous):

y = k/x when y=11 and x=6: 11 = k / 6 k = 66 and equation of variation is y = 66 / x